If North American casino roulette were fair, how much should it pay to a player who bets $10 on red and a red number comes up? The odds against winning a bet on a red in roulette are 1.01 to 1.

Answer by stanbon(57307)

If North American casino roulette were fair, how much should it pay to a player who bets $10 on red and a red number comes up?
---------
The odds against winning a bet on a red in roulette are 1.01 to 1.
---
P(win) = 1/(2.01) ; P(lose) = 1.01/2.01
---------------------------------------------
Note: If the game is "fair" E(x) = 0
Let "x" be casino payback.
Equation:
E(x) = winnings(P(win))+(losings)(P(lose)) = 0
(x-10)(1/2.01)+(-10)(1.01/2.01) = 0
Multiply thru by 2.01 and solve for "x"
---
(x-10) -10(1.01) = 0
x = 10 + 10.1 = $20.10
=============================
Cheers,
Stan H.